# A General Infrastructure and Workflow for Quadrotor Deep Reinforcement Learning and Reality Deployment

### v1

@@ Abstract
Deploying robot learning methods to a quadrotor in unstructured outdoor environments is an exciting task. Quadrotors operating in real-world environments by learning-based methods encounter several challenges: a large amount of simulator generated data required for training, strict demands for real-time processing onboard, and the sim-to-real gap caused by dynamic and noisy conditions. Current works have made a great breakthrough in applying learning-based methods to end-to-end control of quadrotors, but rarely mention the infrastructure system training from scratch and deploying to reality, which makes it difficult to reproduce methods and applications. To bridge this gap, we propose a platform that enables the seamless transfer of end-to-end deep reinforcement learning (DRL) policies. We integrate the training environment, flight dynamics control, DRL algorithms, the MAVROS middleware stack, and hardware into a comprehensive workflow and architecture that enables quadrotors’ policies to be trained from scratch to real-world deployment in several minutes. Our platform provides rich types of environments including hovering, dynamic obstacle avoidance, trajectory tracking, balloon hitting, and planning in unknown environments, as a physical experiment benchmark. Through extensive empirical validation, we demonstrate the efficiency of proposed sim-to-real platform, and robust outdoor flight performance under real-world perturbations. Details can be found from our website https://emnavi.tech/AirGym/.
I Introduction
Using learning-based methods on quadrotors has gained significant attention in the last year. Previous works have demonstrated that end-to-end learning can effectively facilitate the intelligence under complex control tasks, such as drone racing[12, 40, 29, 48, 21], obstacle avoidance[37, 46], trajectory tracking[3, 25, 24], agile flight[45, 48, 8, 44], as well as navigation and planning[46, 48, 16], which presents potential for improved control and cognition performance. In this context, the framework of deep reinforcement learning (DRL) combined with environment interactive learning provides an excellent paradigm for end-to-end quadrotor learning. The integration of sensing enables quadrotors to perform end-to-end inference, significantly shortening the reasoning chain from perception to planning and control, thereby making them more intelligent.
Figure 1: Dynamic obstacle avoidance by end-to-end DRL Sim-to-Real using depth sensing. The quadrotor with blue halo tries to dodge a fast flying football. The experiment is conducted in a wild GNSS-deny environment and only relies on onboard sensing for perception and localization. The speed of flying ball reaches about $15\mathrm{m\cdot s^{-1}}$.
DRL methods and differentiable physics-based approaches have achieved remarkable results on quadrotor agile flight tasks in recent advancements[44, 8, 45, 48, 13, 1, 11, 3, 29, 25]. However, there are still challenges. Most of these works rely on simulation or are achieved with the assistance of indoor or external positioning systems, while only a few of them attempt to transfer policy to the outdoor. The reasons can be primarily attributed to two main aspects. From the control perspective, RL policies trained in simulation are not inherently robust and often exhibit instability when deployed in the real world[3]. Besides, complex environments and disturbances resulting in inaccurate state estimation can cause policies acquired in simulation to fail in real-world scenarios. Without external positioning systems, it is hard to obtain an accurate state estimation of the quadrotor in the wild. Furthermore, devices equipped with quadrotors must achieve onboard inference capabilities, which imposes specific requirements on integration and computation.
Although some studies in training a quadrotor by DRL have provided skills and tricks for reward settings[30, 46, 22, 45, 1, 23, 12, 15, 47, 30, 8, 48, 13, 40], the observation design and reward design have not been summarized and refined, especially from a sim-to-real perspective. Meanwhile, skills for addressing sim-to-real gaps also need a summary. Most importantly, existing quadrotor works based on robot learning methods describe a lot on how to implement training but lack an introduction to underlying tools, and there are few open-source PX4-based toolkits to facilitate the sim-to-real transfer. It is worth mentioning that Agilicious[5] indeed provides an open-source software and hardware system, while it only permits limited open-source use and does not allow further redistribution. Moreover, Agilicious is based on BetaFlight as the low-level controller, which is less used rather than PX4 in the industry. This is also a major constraint on further extending the embodiment methods of aerial robotics to industrial applications.
Figure 2: Proposed sim-to-real platform: X152b. A racing quadrotor is used as the platform and then equipped with the perception device RealSense D430 (https://www.intelrealsense.com/stereo-depth-modules-and-processors/?wapkw=d430) to provide a binocular infrared vision and depth image and an ARM architecture chip powered onboard computer Khadas VIM (https://www.khadas.com/vim).
Figure 3: The pipeline of proposed workflow. The relationship between one hardware platform X152b and four software toolkits including AirGym, rlPx4Controller, AirGym-Real, and control_for_gym are shown in this pipeline. The simulation part and reality part are closed loops and can be independent of each other. It is worth mentioning that the VIO part is not indispensable in our benchmark because global states estimation is not necessary in our Ego-centric design.
Therefore, our motivation is to build a solid infrastructure that facilitates both DRL training and real-world deployment, and conduct a comprehensive study on quadrotor robot learning by using DRL. We attempt to cover general technical aspects of quadrotor DRL in a single work, providing a solid and comprehensive platform for a quick start. The platform enables the efficient end-to-end training tasks that incorporate perception and validate its performance in outdoor scenarios. Although there are already some simulation platforms for quadrotor reinforcement learning, like Aerial Gym[15], OmniDrones[47], CrazySim[20], and Flightmare[38], our work takes it a step further, focusing on rich tasks and quick successful deployment in real world.
In this work, we propose a DRL platform and sim-to-real system built upon existing open-source drone communities. The system integrates the vectorized environment RL simulation platform, the PX4 autopilot system, and visual pose estimation methods. To this end, we developed and open-sourced the following works:
  • Quadrotor platform X152b: an open-sourced onboard sensing and inference flight platform, shown in Fig.2. Where we integrate with powerful computation and perception devices and provide the mechanical model for simulation alignment.
@@ I Introduction
  • DRL simulation platform AirGym [^1] https://github.com/emNavi/AirGym: a large-scale parallel environment simulation based on IsaacGym[27], where we implement five different types of DRL tasks and provide APIs for policy training at various levels of the control hierarchy.
@@
  • Parallel flight controller rlPx4Controller [^2] https://github.com/emNavi/rlPx4Controller: a parallel simulation flight geometric controller implemented in C++ and encapsulated with Python interface, strictly aligned with the logic and parameters of the PX4 autopilot. It flexibly provides control at various levels, including position and yaw ($PY$), linear velocity & yaw ($LV$), collective thrust & attitude angle ($CTA$), and collective thrust & body rate ($CTBR$).

### v2

# Towards Task-Oriented Flying: Framework, Infrastructure, and Principles

@@ Abstract
Deploying robot learning methods to aerial robots in unstructured environments remains both challenging and promising. While recent advances in deep reinforcement learning (DRL) have enabled end-to-end flight control, the field still lacks systematic design guidelines and a unified infrastructure to support reproducible training and real-world deployment. We present a task-oriented framework for end-to-end DRL in quadrotors that integrates design principles for complex task specification and reveals the interdependencies among simulated task definition, training design principles, and physical deployment. Our framework involves software infrastructure, hardware platforms, and open-source firmware to support a full-stack learning infrastructure and workflow. Extensive empirical results demonstrate robust flight and sim-to-real generalization under real-world disturbances. By reducing the entry barrier for deploying learning-based controllers on aerial robots, our work lays a practical foundation for advancing autonomous flight in dynamic and unstructured environments.
1 Introduction
Aerial robots have emerged as revolutionary technologies that drive the development of the low-altitude economy[1, 2]. Their applications span critical domains including emergency rescue operations, precision agriculture, and urban air mobility, with the boosting of global drone market. This explosive growth is fueled by advancements in autonomous capabilities that enable complex operations in unstructured environments. Recent advances in lightweight sensing technologies and edge computing further accelerate their deployment across industries.
Deep reinforcement learning (DRL) has demonstrated superior performance in time-critical tasks including drone racing[3, 4, 5, 6, 7, 8], dynamic obstacle avoidance[9, 10], rapid trajectory tracking[11, 12, 13], agile flight[14, 7, 15, 16], as well as navigation and planning[10, 7, 17], achieving unprecedented perception and decision-making efficiency. Although many studies have provided methods and skills[16, 15, 14, 7, 18, 19, 20, 11, 6, 12] to improve algorithms and give experience on solving specific tasks, the principles of training and deploying a policy by DRL have not been summarized and refined, especially from a comprehensive and thorough perspective covering simulation, deployment, and their bridge. Most importantly, existing quadrotor works based on robot learning methods describe a lot about methodology but lack guidelines, principles, as well as underlying tools for implementation, and there are few open-source toolkits to facilitate the DRL training and transfer. These absences create significant barriers to practical application: researchers often struggle and feel confused about where to begin when designing and deploying a new task-oriented embodied flight task.
Our research aims to consolidate existing end-to-end DRL methods for aerial robots while incorporating our extensive experimentation and trials, distilling general task design techniques, and summarizing training and deployment guidelines for quadrotor DRL, thereby facilitating broader adoption by researchers. Specifically, we refine principles for designing successful quadrotor DRL tasks and consolidate the factors that affect final performance during the sim-to-real process. Additionally, to support our work and facilitate broader research efforts, we propose a fully open-sourced and robust infrastructure and workflow for quadrotor DRL, including simulation environments, software, firmware, and hardware platforms. We also introduce comprehensive experiments across several different types of aerial robot learning tasks. Our proposed platform supports training and sim-to-real transfer in a remarkably short period of time, significantly accelerating algorithmic iteration.
2 Results
Figure 1: A DRL-based sim-to-real task-oriented framework. The left gear serves as the driving gear and represents simulation-related items, including simulator core, task principles middle layer, and sim-to-real techniques outermost layer. The right gear is the driven layer, involving hardware devices as inner layer, as well as firmware and software as outer layer.
2.1 Framework
The task-oriented framework is illustrated in Fig.1. A pair of gears is employed as a metaphor to illustrate the interdependent relationship between simulation-based training and real-world deployment in aerial robot DRL. The left gear functions as the driving gear, symbolizing the training of policies within the simulator, while the smaller right gear serves as the driven gear, representing the deployment of these policies onto quadrotors in the real world.
Simulator is the core of the driving gear, for training a policy. The interactive simulation environment, built upon physics engines and rendering systems, serves as the training arena for DRL. The physics engine simulates and approximates Newton’s physical laws, forming the dynamical foundation for all complex flight behaviors. Concurrently, environmental perception and pattern recognition inherently depend on rendering verisimilitude, where graphical frame-by-frame progression drives the interactive learning process. Built upon the simulator, task-driven policy training requires systematic design considerations. While end-to-end quadrotor missions can encompass diverse task typologies, their design framework must address three fundamental dimensions: Control & Planning, Environmental Sensing, and Training. This might involve multiple aspects, including command hierarchy selection, reward-punishment mechanisms, training environment design, and algorithm adaptation, each involving sophisticated techniques and diverse approaches. At this conceptual level, we systematically abstract and distill seven fundamental design principles addressing the aforementioned three dimensions, relying on both established literature and extensive empirical validation from our own practical works: ① appropriate control modes ② spatiotemporal awareness guidance ③ egocentric sensing and unique observation ④ efficient perception encoding and spatial description ⑤ training curriculum ⑥ random environmental policy ⑦ policy gradient via differential simulation. The detailed explanation of these principles will be comprehensively elaborated later. The outermost layer of driving gear is factors and tricks that directly influence the performance of sim-to-real deployment.
The driven gear can be conceptualized as the standard architecture of an autonomous unmanned system, embodying the following fundamental characteristics: platform, hardware device with firmware, and software application. As the validation platform for this framework and workflow, the self-developed aerial platform X152b is centrally positioned within the gear diagram. Hardware components such as the onboard PC, flight controller, and sensors are integrated into this platform. Sensing by vision[21, 8, 22, 23, 24, 25, 26, 7, 27], LiDAR[26] and neuromorphic vision[28, 3, 29] as well as other perception devices can also be employed. At the outermost layer, the associated firmware & software, state estimation algorithms, and open-sourced communities serve as the medium for transferring and deploying a neural network onto the physical platform, and meshing with the outermost layer of driving gear.
2.2 Infrastructure & Workflow
To facilitate end-to-end policy training and deployment, we integrate insights from numerous existing RL studies along with our practical experience, and propose a general, task-oriented workflow for rapidly training and deploying DRL-based aerial robot flight policies. The pipeline of our platform is illustrated in Fig. 2, where we have implemented two separate closed-loops for training and deployment, tailored for simulation and real-world environments respectively. Our workflow pipeline includes:
@@ 2.2 Infrastructure & Workflow
  • DRL simulation platform AirGym [^1] https://github.com/emNavi/AirGym: a large-scale parallel environment simulation based on IsaacGym[30], where we implement four DRL tasks: tracking, avoidance, target hitting, planning, as well as the most basic flight control task hovering.
@@
  • Parallel flight controller rlPx4Controller [^2] https://github.com/emNavi/rlPx4Controller: a parallel simulation flight geometric controller implemented in C++ and encapsulated with Python interface, strictly aligned with the logic and parameters of the PX4 autopilot. It flexibly provides control at various levels, including position & yaw ($PY$), linear velocity & yaw ($LV$), collective thrust & attitude angle ($CTA$), and collective thrust & body rate ($CTBR$).